UT September 13, 2019 - Restart Night #12


OSA: Josh Williams

ISp: G. Bechetii

Mtn Manager: M. Wagner

Observers: Greg Taylor, Olga Kuhn, John Hill, Barry Rothberg (Tucson)

Summary

Pretty clear skies. The following tasks were done:

  • checkout of SOUL on SX. Confirmed working in all bins. (That's just the barebones summary)
  • AGW1/LUCI1 + AGW2/LUCI2 transform verification data collected.
  • Bino LUCI science-like LS observation using G210
  • Bino MODS science-like LS observation using dual grating mode (PN NGC40).
  • Bino MODS check on blind offsets in Stone A field using 5" slit. By-eye looks good.
  • Bino MODS check on NS guiding with dithering.
    • No probe timeout errors (IT 7845). WFS exptime manually dropped to 5s and stuck, but star was much brighter than the previous night when the set time was overridden by dynamic scaling (IT 7643).
  • LBC non-sidereal observation to check guiding & dithering.
    • No guiding failures. However, exptime was 3x shorter (60s vs 180s) than when we experienced guiding failures. Maybe this is a clue (IT 7847)

AO work

At sunset, Greg plans to start with AO checkout.

UT 02:20 - proceeding with AO SOUL on-sky checks, awaiting AGW transforms

UT 02:39 - configuring LUCI-1 for N30 fieldstop, FeII +clear, cube readout

L1 0001-0002 - 2.51 sec test dark

L1 0003 -0004 test exposure on AO target (accidentlly took two exposures)

-too bright, going to H~7 target

L1 0005 - test exposure - 2.51sec

-core saturated, but G. Taylor still going to use it

L1 0006 - test exposure 2.51sec x 5 NDIT cube

UT 03:10 - moving to new target, Bin 1 test successful

UT 03:25 - D. Miller has provided a new set of transforms for LFBG and RFBG, J. Hill will update. We will alsup update PCS config for RFBG zp rotator change

L1 0013 -2.51 sec test looks good

L1 0014 - 2.51sec x 5 ndit cube - looks good

UT 04:19 - updated PCS config on TCS
  • Correct path is: /lbt/data/config/tcs/PCS on tcs1. In that PCSInstrument.conf file, RIGHTZEROPOINT was 177.05. Copied the conf file used on 20190907 to PCSInsrument.conf, set RIGHTZEROPOINT to 177.30.
  • will need to restart PCS and GCS
UT 04:27 - putting N30 fieldstop in LUCI-2, first rotated to 210
  • LUCI-2 MOS error, El 78, Rotator 210
  • Attemptd MOS storage to FPU
    • MHU translator to mask OK
    • MHU rotator OK
    • selcted mask
    • MHU rotator to mask OK
    • Failed at MHU picker arms grab mask - zeroed strain gauges - going for mask slot 2
    • Masks are still locked
      • grabber did 8000 steps, then 500 steps, then 1000 steps
      • Slot 2 position normally is -1352 translator position and angle 266.5
      • Current position is -1353 angle=266.04
      • storage angle tolerance = 0.5
    • not limits are on grabber - neg would be not grabbed
    • strain gauges are on
    • Recovery
      • setting 1.1 steps to 0
      • moving -8000 steps
      • negative limit is on
      • moved -4734 steps
      • initializing
    • trying again
    • Worked! N30 Fieldstop in FPU
UT 05:25 - G. Taylor done with AO checks

UT 05:27 - restarting PCS and both GCS to absorb new transform and LUCI-2 rotator ZP

-authorizing for binocular LUCIs, setting up LUCI-1 for AGW transform

UT 05:30 - P. Kubanek will set up MCSPU update, J. Wiliams and G. Taylor resting shells

UT 05:33 - also seeting elevation limit to 25degrees to view comet later

-will go to Stone O to point and collimate, then used modified list of 12 stsrs "Stone_O_m12.list" PA=170

UT 05:41 - sending preset to Stone O

UT 05:45 - spiral search for pointing star - didn't land close to it - also 20 degrees from Moon, star is R=11.6

-going to an M5 first, then will return to Stone O to collimate

UT 05:59 - collimating on Stone O 0 - exp=4.0 H-band N30 fildstop, N3.75 camera
  • Stopping Scripts both sides EXP=4.0 too bright
-stooping DX side its taking 2 exposures at a time, setting NEXP=1 and taking 2.51 sec waste expsure (0007)

-ignore L1 0039-0044 and L2 0003-0010

Settiing to DIT=2.51, NDIT=1 H-band

L1 0045 L2 0011 Star 0

L2 0012 Blank

Still too bright need to switch FeII filters

L1 0046 test image - counts look good

UT 06;19 - trying this again with FeII 2.51sec
  • L1 L2
  • 47 Blank 13
  • 48 Star 11 14
  • 49 Star 15 15
  • 50 Star 24 16
  • 51 Star 23 17
  • 52 Star 28 18
  • 53 Star 26 -LUCI-2 didn't collect thinks it should be #0018
  • 54 Star 33 19
  • 55 Star 32 20
  • 56 Star 12 21
  • 57 Star 10 22
  • 58 Star 14 23
  • 59 Star 1 24 -vignetted by guide probe - bad choice by me
  • 60 Star 0 25

Going to run one more list "Stone_O_p10.list"
  • 61 Star 0 26
  • --LUCI-1 did not take an image it requested one but called it #0061 blank 27
  • 62 Star 11 28
  • 63 Star 15 29
  • 64 Star 24 30
  • 65 Star 23 31
  • 66 Star 28 32
  • 67 Star 26 33
  • 68 Star 33 34
  • 69 Star 32 35
  • 70 Star 12 36
  • 71 Star 10 37
  • 72 Star 0 38
UT 07:05 - DX guide star jumped

UT 07:06 - completed 2nd set of transform

UT 07:09 - reversed PCSInstrument.conf for RFBG ZP to 177.05 - upon confirmation new value works change back to 177.30 AND DO THIS FOR ARGOS TOO

Going to run test script to check on vignetting issue IT 7235

UT 07:14 - running Stone 0 dither PA=0

-pointing on right side failed and its waiting for collimation on left

this time WFS went through on SX left rotator = 331

-sending PA=180 - pointing missed again on DX, SX found OK LFBG rotator = 151

-sending PA=270 - pointing missed again on DX

-sending PA=90 - pointing missed again on DX
LFBG
07:18 AGW position (mm): -139.443, 83.0191 (native: 140.013, 529.522)
rot=331

07:19  AGW position (mm): -142.383, 65.9863 (native: 142.943, 546.586)
rot=151

07:23 AGW position (mm): -137.659, 69.2416 (native: 138.212, 543.315)
rot=240

07:27 AGW position (mm): -140.989, 79.7684 (native: 141.559, 532.781)
rot=419

The following are extracted from FITS headers for the above experiment (and a few on-axis positions beforehand).
gethead left_wfscimage0003*.fits TIME-OBS ROTANGLE PROBEX PROBEY AGWX AGWY
Beware of moonlit clouds and changing background levels.
FILENAME                  TIME-OBS     ROTANGLE         PROBEX          PROBEY         AGWX    AGWY
left_wfscimage000328.fits 06:48:15.148 172.84790039062 0.052457098001556 -0.20668958281964 0.161 612.536
left_wfscimage000329.fits 06:52:57.696 171.40927124023 0.058685039040205 -0.20908620885228 0.147 612.539
left_wfscimage000330.fits 06:58:38.945 169.72831726074 0.066331022206015 -0.21209511197175 0.161 612.541
left_wfscimage000331.fits 07:03:21.582 168.38410949707 0.072828423753822 -0.21479102658907 0.132 612.544
left_wfscimage000332.fits 07:05:43.915 167.72476196289 0.076144599875566 -0.2162293477582 0.117 612.546

left_wfscimage000333.fits 07:18:22.745 331.44332885742 -139.45117229704 83.02187811129 140.039 529.515
left_wfscimage000334.fits 07:18:48.740 331.34436035156 -139.45117229704 83.02187811129 140.039 529.515
left_wfscimage000335.fits 07:19:14.509 331.24536132812 -139.45117229704 83.02187811129 140.039 529.515

left_wfscimage000336.fits 07:21:30.745 150.79907226562 -142.23924245864 65.947982046093 142.943 546.586
left_wfscimage000337.fits 07:21:46.717 150.73968505859 -142.23924245864 65.947982046093 142.943 546.586
left_wfscimage000338.fits 07:22:12.418 150.64273071289 -142.23924245864 65.947982046093 142.943 546.586
left_wfscimage000339.fits 07:22:38.575 150.5453338623 -142.23924245864 65.947982046093 142.943 546.586
left_wfscimage000340.fits 07:23:07.137 150.43910217285 -142.23924245864 65.947982046093 142.943 546.586

left_wfscimage000341.fits 07:24:40.040 240.13238525391 -137.69730599592 69.251090070806 138.186 543.325
left_wfscimage000342.fits 07:25:08.177 240.0304107666 -137.69730599592 69.251090070806 138.186 543.325
left_wfscimage000343.fits 07:25:33.973 239.93574523926 -137.69730599592 69.251090070806 138.186 543.325
left_wfscimage000344.fits 07:25:59.814 239.84150695801 -137.69730599592 69.251090070806 138.186 543.325
left_wfscimage000345.fits 07:26:25.798 239.74862670898 -137.69730599592 69.251090070806 138.186 543.325

left_wfscimage000346.fits 07:28:27.967 419.16555786133 -140.96849735269 79.761599444731 141.571 532.774
left_wfscimage000347.fits 07:28:43.926 419.10916137695 -140.96849735269 79.761599444731 141.571 532.774
left_wfscimage000348.fits 07:29:10.037 419.01742553711 -140.96849735269 79.761599444731 141.571 532.774
left_wfscimage000349.fits 07:29:35.863 418.92608642578 -140.96849735269 79.761599444731 141.571 532.774
left_wfscimage000350.fits 07:30:01.820 418.83560180664 -140.96849735269 79.761599444731 141.571 532.774
left_wfscimage000351.fits 07:30:27.701 418.74526977539 -140.96849735269 79.761599444731 141.571 532.774

UT 07:30 - doing a pointing check before G210 test scripts

-missed on both sides - ever since yesterday pointing has been attrocious

UT 07:36 - sending preset to Telluric Star HD210290

-center in slit at X=1030

L1 0073-0077

L2 0039-0043

-factor of 3 difference beween L1 and L2, but I thought it was well centered

UT 07:49 - going to QSO target 4C+31.63

LUCI-2 MOS error EL 65, RFBG rotator 283
  • FPU to turnout
  • Grab mask failed
  • again, 8000 moves, then 500 steps, then 1000 steps on grabber from log
  • strain gauges are on, Maek 7 (0.5" slit is open)
  • picker arms are open, FPU is hold but not closed
  • Recovery
    • 1.1 grabber , set position to 0, moving grabber -8000, steps moved was -5546
    • Next step move backwards - FPU hold to lock, pre-conditions say true at hold, verified with switch box that FPU is hold and clamp x and clamp y has reference on (1.3,1.4)
    • executing hold to lock fpu
    • switchbox shows FPU closed and 1.3 and 1.4 neg limit on
    • initialize, then check holding currents, turn off rotator HC, initializatin turned HC off in spindle 1.5
Recovered at UT 08:09

-puttig in slit at 1030, offsets +1.58", -2.66" L1, -0.42" -1.6" L2, adding +0.12" DETX

taking 2.51x12 confirmatory image

UT 08:24 - starting spectra

L1 0078 -0081 acq

L2 0044-0059 acq (the extra 12 were because I took 12ndit but saved as normal)

-definitely clouds in sky, WFS 0.8-0.9" both sides

Spectra:
L1 0082-0086

L2 0060-0064

UT 08:53 - done with LUCIs on-sky tonight

-putting LUCIs in safe config until cals later

UT 09 Switching instruments from bino LUCI to bino MODS. Problems authorizing "Authorize busy". Josh tried a few times, but without avail. DG focal stations showing up on some TCS GUIS, but getting errors. Calling Petr. Problem resolved.

09:39 Taking a long time to collimate on SX, but DX is fine. Tried a brighter star and cleared active optics. Still no joy. Homing the hexapod again. That was done at the start of the night. MODS1 AGW was homed before we started. AGW5 - RPC error from the OAC -- but meaningless as MODS doesn't comm through OAC.

09:54 Looking better. Apparently homing hexapod did the trick.

10:00 on NGC40.
  • Slit image m1r 2 and m2r 2. 10s
  • Field image m?r 3. 10s - sat. Another with 2s 4
  • Confimatory thru-slit image 5. Tweaking mods2 by just -0.1", while mods1r is still reading out. Tweaking mods1 by -0.24".
  • Resulting pair of images, 6, look pretty good. Seeing FWHM 0.9".
  • Spectra: m1b 2--4, m1r 7-9 is , m2b 2-4, m2r 7-9

10:17 Blind Offset Test on Stone A.
  • acqBinoMODS StoneA_23to17.txt (preset completed but I control-c'ed script running windows, little time left.)
  • acqMODS --mods1 -a StoneA_23to17.txt
  • acqMODS --mods2 -a StoneA_23to17.txt
  • forgot the bino flag...
  • repeat whole acqBinoMODS
    • MODS1 Offset Command: offsetxy -0.367 11.089 rel
    • MODS2 Offset Command: offsetxy 4.053 7.125 rel
  • Images m1r 13 and m2r 13 are the confirmatory through slit images.
  • DONE: BC Sent BinocularControl command SYNCOFFSET, M1> OFFSET -116.430 119.620 abs
  • and m1r 14 and m2r 14 are the images after the blind offset.

10:47 MODS guiding and dither
  • 2018 YH.
  • Set WFS exptime of 5-sec. WFS time is sticking at 5-sec today. This time spots are pretty bright (star has R=14.9).
  • Not seeing any timeout errors.
  • Slewing to change config while reading out... let's see how this does.

11:05 Reconfiguring to the LBCs.
  • 11:25 Reconfigured and 1st preset completed.
  • 11:25 running dohybrid a bit higher than comet.
    • dohybrid detected and sent coma, ~2000nm, which looks correct.
    • blue converged in 2 FPIA + 1 WRS iteration and red ... shift-S
  • 11:34 go to CP field for comet. Airmass 2.367 elev 25 deg (well, no... it was 24.x deg).
    • adsecs ripped. have to go to 40 deg to set them.
  • 11:41 back in business. Sending preset to Comet CP field.
    • running dofpia. Estimated seeing is high, and blue converged already.
    • 11:47 converged in 3 iterations. Estimated seeing is 1.6 on blue and 1.3 on red, so that helps.
  • 11:48 copointing images.
    • b 114820 and r 114816
    • offsets were (hoping I selected the same star on both sides - a companion was hard to see on the blue):
On LBC-Blue/LBC-Red the star is 23.9/8.9 arcsec away from the rotator center
Sending the calculated optics offsets:
SX: dX = 0.171 dY = 0.344 dRX = 7.375 dRY = -3.668
DX: dX = -0.200 dY = -0.357 dRX = -7.655 dRY = 4.288

  • 11:52 starting observation and guiding. Observation will take 5 60sec exposures dithers at g and r
    • Guiding looks ok --- so far, so good.
    • We can see comet - it is pretty faint, but the sky background is high, 30k at g and 20k at r.
    • 120419(b) and 120414(r) pausing to enter a scaling factor of 0.5 because background at g was rising to 35k. This concludes our guiding test (because 30 is less than the 40-sec threshold).
      • We saw no "no star found" errors this time. John said this error arises when the guide star is a certain distance from its reference position, measured on the first guider image. This was put in to
  • 12:10 Repeated script. Forgot to click set. 121024(b) 121014(r) are sidereal. Clicked set but LBC is still taking exposure. Play buttom enabled. Odd... LBC is continuing to take exposures... I cannot stop it! And if I click the Play button, it has no effect. 12:12 UT.
  • 12:18 Josh cancelled the preset, but LBC is still going. Slew up to zen, away from E.
  • First set of 9 dithers is non-sidereal, 2nd set is sidereal.

  • 12:23 Skyflats at V and r. I realized I should have started with g and i. So I stopped the OB before the slew completed and uploaded the g and i one. The test skyflat was taken before the telescope was on-target. b 122655 and r 122653.
  • 12:29 g and i flats preceded by another (accidental) test.
  • then g and R flats @ PA 180
  • then B and r flats @ PA 180
  • then just one S and F @ 180
  • one U and F @ 180, as U is saturated, continuing with F only. I had clicked pause to turn off blue. When OB finished, counter read 8/10.
  • now F @ 180. scale was 0.5 paused while reading out.

MODS final tests in non-phot skies:
  • Observing a PN through 0.6" slit in dual grating and dual prism modes.
  • Check Blind Offsets on StoneA
  • Check NS + guiding and dithering.


Plan Outline

  • DONE LUCI-Bino AGW Transform confirmation - Highest Priority
  • LUCI-Bino G210 science-like observations
  • DONE AO-Onsky checkout
    • SX/FLAO/SOUL on-sky check-outs. (high)
    • Bin1 1, 2, & 4 and NCPA verification.
  • Test patch to GCS/PCS to address dithering NS guiding
    • LUCI
    • MODS
  • DX M3 alignment
  • LBC
    • Standard photometric fields - requires clear skies (any seeing) - required for ZP calibrations
    • Spectro-photometric standards - requires clear skies (any seeing) - required for ZP calibrations
    • Twilight Flats all filters
    • LBC sub-framing (test OBs needed)
    • NS guiding with dithers
    • Checkout observations with Bessel filters
  • MODS
    • Test NS guiding with dithering

-- BarryRothberg - 12 Sep 2019
Topic revision: r9 - 17 Sep 2019, JohnHill
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback